[Ncep.nhc.nco_contacts] UPDATE: TSB Morning Rounds - Mon Aug 28, 2017

Pablo Santos pablo.santos at noaa.gov
Mon Aug 28 12:56:36 UTC 2017

I added the following item as well. Pablo

1.) PETSS did not run for the 00Z and 06Z cycles. Pablo brought it up at 
the 8 AM NCO and Doug F was going to follow up on this.

On 8/28/17 8:52 AM, TSB Admin wrote:
> Here is today's summary of NHC computer operations:
> --- Pablo
> ----------
> *TAFB:*
> 1.) The Cable TV remains inoperable on the upper right west wall 
> situational awareness display in TAFB and the large center TV on the 
> east wall. NCO ticket open. Comcast coming in today to take care of this.
> 2.) Large monitor on EPAC desk, LX3 has a faulty video connector, and 
> can cause video display issues is not treated tenderly. Doug G taking 
> a look at this.
> 3.) The automated danger graphic shows "CYCLONE" for Potential 
> Tropical Cyclones. Is this correct? Matt Sardi to look into it.
> 4.) Security Camera display in WFO Miami and TAFB ops areas is frozen 
> to a time from yesterday during daylight with "video loss" red banner 
> showing on all 10 cameras. A new NCO TT was opened.
> 5.) AC Cooling Unit #3 in the server room had an alarm beeping last 
> night and again tonight when we came in. Alarm display said "Short 
> Cycle Comp.2". Cleared the alarm by pressing Enter. No issues noted 
> with the temperature on the display - WIlliam Notified.
> *HSU:*
> 1.) The 21Z advisory graphics did not launch on 8/26. Chris Mello was 
> able to troubleshoot the issue and get them released. If any issues 
> remain, Chris to work with Matt Sardi on this.
> 2.) On 8/26 it was reported that the in-house JHT genesis tool 
> continues to not function correctly, as the UKMET output is not being 
> included both on its own as well as in the consensus. Craig Mattock to 
> investigate.
> 3.) The SHIPS guidance for Harvey the 26th at 00Z did not print. 
> Looking on the ATCF ftp site, the file for Harvey was available but it 
> was an empty file. We had to resubmit guidance for Harvey and AL92 
> (twice) as the SDM could not see the initial submission. Issued 
> appears to have been resolved.
> *Other items:**NHC RFCs:*
> 1.) RFC 3674 to allow wpse account access to opah scheduled for 
> Wednesday this week.
> 2.) RFC 3675 - Add NHC printer to NCEP Mail Relay - scheduled for 
> wednesday this week. This change will allow a printer (, 
> or carmenrohp.nhc.noaa.gov) to be able to scan to email with 
> mrelay.ncep.noaa.gov port 25. This worked until recently.
> 3.) RFC 3647 - NHC Netapp snapshot policy changes scheduled for Wed 
> weather permitting.
> 4.) RFC 3651 - NHC Fix for Single Point DNS and NTP Servers scheduled 
> for Thursday.
> 5.) RFC 250 - Automated "Key Messages Graphic" - postponed due to 
> Harvey. But will be taken care of as soon as possible.
> 6.) RFC 251 - Add email alert for satellite purge job failure. 
> Proposed to implement on 8/30
> 7.) RFC 252 - Upgrade ATCF FTP scripts to rsync. Proposed to implement 
> on 8/30
> 8.) RFC to hide gTCM files in NHC ftp site still going being written.


Dr. Pablo Santos
Acting Chief
Technology and Science Branch
National Hurricane Center, Miami, FL
(c) 954-559-4965
Facebook: https://www.facebook.com/NWSNHC/
Twitter: @NWSNHC @NHC_Atlantic @NHC_Pacific  @NHC_TAFB  @NHC_Surge

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170828/af9ed041/attachment-0001.html 

More information about the Ncep.nhc.nco_contacts mailing list